OpenStudy (anonymous):

Find the exact circumference, in terms of , of a circle with the given radius. 5 inches

OpenStudy (anonymous):

C=2(3.14)r C=2(3.14)(5)

OpenStudy (anonymous):

Circumference=2*r*pi Sooo, circumference=2*5*pi = 10 * pi More simply 10pi